OpenStudy (rissa):

Solve the proportion for the given variable. 24/4 = 18/b

OpenStudy (anonymous):

b = 3

OpenStudy (agreene):

24/4 = 18/b 6 = 18/b multiply both sides by b/6 b= 3

OpenStudy (eyust707):

funny how everyone thinks of math differently

OpenStudy (eyust707):

i would have cross multiplied